OpenStudy (anonymous):

∛(√(3x+1)) how do you simplify this?

OpenStudy (asnaseer):

do you know how to write square root as an index, i.e.:\[\sqrt{x}=x^{?}\]

OpenStudy (anonymous):

x^1/2

OpenStudy (asnaseer):

correct, similarly, do you know how to write cube roots in index form?

OpenStudy (anonymous):

x^1/3 ?

OpenStudy (asnaseer):

perfect, so you can rewrite your expression as:\[\sqrt[3]{\sqrt{3x+1}}=\sqrt[3]{(3x+1)^{\frac{1}{2}}}=((3x+1)^{\frac{1}{2}})^{\frac{1}{3}}\]and then use the law of indexs to simplify this

OpenStudy (anonymous):

thank you!!

OpenStudy (asnaseer):

yw :)